You are here

Beam delivery update

header

09 December 2014

Today’s planned beam development is still underway but is scheduled to wrap up by 20:30. BL1A has been defined off for the year; high-energy beam will be delivered to BL1B for PIF once full cyclotron operation is restored. BL2A is currently online so that a new beam position monitor can be tested, though it will be defined off overnight. Finally, BL2C has been switched from 2C1 (PIF) to STF (2C4) while PIF users run on BL1B; that’s expected to be the case for the next week.

BL2A will be turned off overnight as we have no need to put protons on the ISAC production target – last night’s 213Fr implantation for experiment L112 was successful and we’ll set up beam for FrPNC experiment S1218 tomorrow morning. The target is also midway through its life and the integrated proton current on target looks as though it will end up very close to the 5,000 µA•hr limit for UCx. RIB delivery to S1218 will happen some time tomorrow. Meanwhile, DRAGON will continue to take 76Se from the OLIS multicharge source for detector tests as a problem with the DTL RF put an end to stable-beam delivery at ~15:30 this afternoon. The RF group hasn’t yet been able to fix that, but will continue their efforts tomorrow morning.

-- ACM